OpenStudy (anonymous):

Solve S = 4v2 for v. What steps do you use to solve? I think that you divide 4 from both sides and then find the root? I am lost!

OpenStudy (anonymous):

s=4v^2 √s=√4v √s=2v 1/2√s=v?

OpenStudy (anonymous):

\[s=4v^2\] divide both sides by 4 \[v^2=\frac{s}{4}\] \[v=\sqrt{\frac{s}{4}}\] \[v=\frac{\sqrt{s}}{2}\]

OpenStudy (anonymous):

I get it! Thank you!!!